Find the angle between the lines of intersection of the cone 4π‘₯2 + 𝑦2 + 4𝑧2 + 4𝑦𝑧 + 2𝑧π‘₯ = 0 and the plane π‘₯ + 2𝑦 + 3𝑧 = 0.


1
Expert's answer
2022-02-09T13:38:19-0500

For the first find an equation of intersection curve.

"\begin{cases}
 4x^2+y^2+4z^2+4yz+2zx=0 \\
 x+2y+3z=0 
\end{cases}"



"\begin{cases}
 z=-\frac{x+2y}{3} \\
 4x^2+y^2+4(\frac{x+2y}{3})^2-\frac{4y(x+2y)}{3}-\frac{2x(x+2y)}{3}=0 
\end{cases}"



"34x^2+y^2-8xy=0"

Transform the equation



"(y-(4-3\sqrt{2}i)x)(y-(4+3\sqrt{2}i)x)=0"

This is equation of two imaginary intersecting lines.


Angle between the lines y=k1x and y=k2x is "tan \alpha =\frac{k_{2}-k_{1}}{1+k_{2}k_{1}}"


In our case "k_{1}=4-3\sqrt{2}i" and "k_{2}=4+3\sqrt{2}i" .

It mean, that "tan\alpha =\frac{(4+3\sqrt{2}i)-(4-3\sqrt{2}i)}{1+(4-3\sqrt{2}i)(4+3\sqrt{2}i)}=\frac{6\sqrt{2}i}{34}=\frac{3\sqrt{2}i}{17}"


The angle between the lines of intersection is

"\alpha=tan^{-1}(\frac{3\sqrt{2}i}{17})"